OpenStudy (anonymous):

solve and find the complex roots of 125x^3 + 343 = 0

OpenStudy (anonymous):

sum of two cubes \[a^3+b^3=(a+b)(a^2-ab+b^2)\] in this case you have \[(5x)^3+7^3\]

OpenStudy (anonymous):

factor as \[(5x+7)(25x^2-35x+49)\]

OpenStudy (anonymous):

one root is \[-\frac{7}{5}\] other two complex roots via the quadratic formula

OpenStudy (anonymous):

so I use the quad formula for the 25x^2 - 35x + 49 right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

\[x= \frac{-b \pm \sqrt{b^2-4ac}}{2a}\]